### v2

# To Study Properties of a Known Procedure in Adaptive Sequential Sampling Design

Abstract
We consider the procedure proposed by Bhandari et al., (2009) in the context of two-treatment clinical trials, with the objective of minimizing the applications of the less effective drug to the least number of patients. Our focus is on an adaptive sequential procedure that is both simple and intuitive. Our findings show that the number of applications of the less effective drug is a finite random variable whose all moments are also finite. In contrast, Bhandari et al., (2009) observed that this number increases logarithmically with the total sample size. We attribute this discrepancy to differences in their choice of starting sample size and the method of analysis employed.
Keywords: Adaptive allocation, Average sample number, Composite hypothesis, Incorrect inference probability, Less effective drug application count
1 Introduction
The field of adaptive sequential design has a long history of research, particularly in optimizing multiple objective functions. Berry and Fristedt, (1985) in their book on Bandit Problems, explored Bayesian methods to achieve optimal designs in this context.
In conventional sequential analysis for the two-population problem, samples are drawn one by one at each stage from both populations, ensuring that the sample sizes remain equal. The primary objectives in these cases are to make inferences about the population parameters with less error while minimizing the average sample number.
In contrast, adaptive sampling allows for unequal sampling from the two populations at each stage (sometimes selecting a sample exactly from one population). The choice of which population to sample at each stage depends on the performance of previously collected data. This approach has a natural connection with the context of most studied clinical trials, where two treatments are applied sequentially to a series of patients. The goal in such trials is to balance the need for statistical accuracy with ethical considerations, specifically applying the less effective treatment to a minimum number of patients.
In this context, Bhandari et al., (2007) considered the case with known parameters and concluded that the expected number of applications of the less effective drug is finite under an adaptive sequential design. However, where the parameters are unknown, Bhandari et al., (2009) found that the expected number of applications of the less effective drug increases logarithmically with the total sample size. This result was derived under the assumption of a large initial sample size.
In the present work, we modify the procedure proposed in Bhandari et al., (2009) and provide a more refined analysis to elucidate that the expected number of applications of the less effective drug remains finite. In fact, we show that all the moments of the random variable that denote the number of applications of the less effective drug, are finite.
The paper is structured as follows. We explicate the procedure in Section 2, followed by a theorem with some remarks in Section 3. The proof of the theorem is provided in Section 4. In Section 5, we assess the performance of the procedure under various parameter choices using simulations. Finally, Section 6 concludes.
2 Preliminaries
Let $\mathcal{X}=X_{1},X_{2},X_{3},\ldots$ follow the i.i.d. $\operatorname{N}(\theta_{0},\sigma_{0}^{2})$ and $\mathcal{Y}=Y_{1},Y_{2},Y_{3},\ldots$ follow the i.i.d. $\operatorname{N}(\theta_{1},\sigma_{1}^{2})$ and they be two independent data streams. We will use population 00 and population $1$ for $\mathcal{X}$ and $\mathcal{Y}$, respectively. We draw samples adaptively, i.e., we draw samples sequentially and at stage $n$, after drawing a total of $n$ samples, we define the following two count variables:
$N^{\prime}_{0,n}:=\text{ number of samples drawn from }\mathcal{X},$
$N^{\prime}_{1,n}:=\text{ number of samples drawn from }\mathcal{Y}.$
Let $\hat{\theta}_{0,n}$ and $\hat{\theta}_{1,n}$ denote the sample means at stage $n$ for $\mathcal{X}$ and $\mathcal{Y}$, respectively. We adopt the following allocation rule:
  1. (i)
    If $\hat{\theta}_{0,n}-\hat{\theta}_{1,n}>0$, we increase $N^{\prime}_{0,n}$ by 1,
  2. (ii)
    If $\hat{\theta}_{0,n}-\hat{\theta}_{1,n}<0$, we increase $N^{\prime}_{1,n}$ by 1,
  3. (iii)
    If $\hat{\theta}_{0,n}-\hat{\theta}_{1,n}=0$, we increase either $N^{\prime}_{0,n}$ or $N^{\prime}_{1,n}$ by 1 with probability $\frac{1}{2}$ each.
There are $N$ patients in total. Finally, when $n=N$, we accept the null $H_{0}:\theta_{0}>\theta_{1}$, with probability $1$, if $\hat{\theta}_{0,N}-\hat{\theta}_{1,N}>0$ and with probability $\frac{1}{2}$, if $\hat{\theta}_{0,N}-\hat{\theta}_{1,N}=0$.
Remark 1.
The procedure described above was extensively studied by Bhandari et al., (2009), where it was referred to as Procedure III.
Remark 2.
In the proof of the main result, we assume $\sigma_{0}^{2}$ and $\sigma_{1}^{2}$ to be unity and $\theta_{0}>\theta_{1}$ without any loss of generality.
3 Main Result
Let us define the following:
$N_{1,n}:=\min\{N^{\prime}_{1,n},N^{\prime}_{0,n}\}\quad\text{and}\quad N_{1,N}:=\min\{N^{\prime}_{1,N},N^{\prime}_{0,N}\}.$
Theorem 3.1.
We have the following:
  1. (i)
    $\lim\limits_{N\to\infty}N_{1,N}<\infty$, i.e., $N_{1,N}$ is a finite random variable.
  2. (ii)
    All the moments of $N_{1,N}$ are bounded (uniformly over $N$).
Remark 3.
Bhandari et al., (2009) found that both $\frac{N^{\prime}_{1,N}}{\log(N)}$ and $\frac{\mathbb{E}(N^{\prime}_{1,N})}{\log(N)}$ tend to positive constants, contrary to our main result. This variation arises because they started with a large initial sample size and employed a different method of (mathematical) analysis.
Remark 4.
Note that, $\lim\limits_{N\to\infty}N_{1,N}\neq\lim\limits_{N\to\infty}N^{\prime}_{1,N}$ if and only if there is incorrect inference about $H_{0}:\theta_{0}>\theta_{1}$. The probability of incorrect inference is negligible and it is less than or equal to
$\Phi\left(-\left(\theta_{0}-\theta_{1}+\varepsilon\right)\cdot\sqrt{N_{1,N}}\right)<\Phi\left(-\left(\theta_{0}-\theta_{1}+\varepsilon\right)\cdot\sqrt{M}\right),$
for some $\varepsilon>0$, where $M$ is a finite natural number and $\Phi(\cdot)$ is the CDF of the standard normal distribution. We start with an initial sample size of $M$ from each population at the beginning of the procedure.
Remark 5.
Note that, in the case of sequential two-population testing problem, when both $\theta_{0}$ and $\theta_{1}$ are unknown, it is not possible to achieve consistent hypothesis testing (i.e., PCS approaching 1 as the total sample size tends to infinity) if the expected number of samples from one population remains finite while the other grows indefinitely.
4 Proof of Main Result
The following two results, as stated in Chapter 1, Section 9 of Billingsley, (1995), are presented here without proof.
Result 4.1.
(Law of Iterated Logarithm) Let $\left\{X_{n}\right\}$ be i.i.d. random variables with mean zero and unit variance. Let $S_{n}=X_{1}+X_{2}+X_{3}+\ldots+X_{n}$. Then,
$\limsup\limits_{n\to\infty}\frac{\left|S_{n}\right|}{\sqrt{2n\log\operatorname{logn}}}=1\quad\text{almost surely.}$.
Result 4.2.
(Chernoff’s Theorem in the context of Normal Distribution) Let $\left\{X_{n}\right\}$ be the i.i.d. random variables following $\operatorname{N}(\mu,1)$ with $\mu<0$, with mgf $M(t)$. Suppose $\bar{X}_{n}=\frac{1}{n}\left(X_{1}+X_{2}+X_{3}+\ldots+X_{n}\right)$. If $\rho=\min\limits_{t}M(t)$, then $\rho<1$ and
$\lim\limits_{n\to\infty}\frac{1}{n}\log\left[\mathbb{P}\left(\bar{X}_{n}\geq 0\right)\right]=\log(\rho),$
which implies $\mathbb{P}\left(\bar{X}_{n}>0\right)<\mathrm{C}\cdot\rho^{n}\hskip 5.69054pt\forall\hskip 5.69054ptn$, for some positive constant $\mathrm{C}$.
Remark 6.
Note that, along with normal distribution for $\{X_{n}\}$, Chernoff’s theorem also holds for any distribution with $\mathbb{E}(X_{n})<0$ and $\mathbb{P}(X_{n}>0)>0$ (see Billingsley, (1995)).
Using Result 4.1, we find that for any given $\varepsilon>0$, there exists a natural number $N_{\varepsilon}^{*}$ and subset of sample space $A$ with $\mathbb{P}(A)>1-\varepsilon$, such that
$\forall\hskip 5.69054pt\omega\in A,\quad|\hat{\theta}_{0,(n)}(\omega)-\theta_{0}|<\varepsilon\text{ and }|\hat{\theta}_{1,(n)}(\omega)-\theta_{1}|<\varepsilon\text{, (for all }n>N_{\varepsilon}^{*}\text{),}$,
where $\hat{\theta}_{0,(n)}$ and $\hat{\theta}_{1,(n)}$ are the mean of first $n$ observations coming from the data streams $\mathcal{X}$ and $\mathcal{Y}$ respectively.
Using the above, we find that with probability $\geq(1-\varepsilon)$,
$\lim\limits_{N\to\infty}N_{1,N}<N_{\varepsilon}^{*}\text{, i.e., }N_{1,N}\text{ is a finite random variable.}$.
This proves part (i) of Theorem 3.1.
Let us define
$M_{u}^{*}=\inf\left\{k:\sup\limits_{n\geq k}\bar{Z}_{n}\leq u\right\},$
where $Z_{1},Z_{2},Z_{3},\ldots,Z_{n}$ are i.i.d. $\operatorname{N}(0,1)$ and $\bar{Z}_{n}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$.
Define $X_{i}=Z_{i}-u$ for small $u>0$, for every natural number $i$.
Then, Result 4.2 holds for this sequence of $X_{i}$.
Remark 7.
Let $u=\frac{\theta_{0}-\theta_{1}}{3}$. Then, $N_{1,n}\leq M_{u0}^{*}+M_{u1}^{*}$, for some $M_{ui}^{*}$ ($i=0,1$) which are i.i.d. $M_{u}^{*}$. This is because there exists $M_{u0}^{*}$ and $M_{u1}^{*}$ (independent) such that $\hat{\theta}_{0,(n)}\in(\theta_{0}-u,\theta_{0}+u)$ for all $n\geq M_{u0}^{*}$, $\hat{\theta}_{1,(n)}\in(\theta_{1}-u,\theta_{1}+u)$ for all $n\geq M_{u1}^{*}$ and $N_{1,n}=min\{N_{1,n}^{\prime},N_{0,n}^{\prime}\}\leq N_{0,n}^{\prime}+N_{1,n}^{\prime}\leq M_{u0}^{*}+M_{u1}^{*}$.
The proof of part (ii) of Theorem 3.1 will be complete upon proving Lemma 4.1.
Lemma 4.1.
The mgf of the random variable $M_{u}^{*}$ is finite in a small open neighbourhood of 0 i.e., for $t\in(-\delta,\delta)$ where $\delta>0$ and $\delta$ is small, $\mathbb{E}(\exp(tM_{u}^{*}))<\infty$. Hence, all the moments of the random variable $M_{u}^{*}$ are finite.
Proof.
We will prove the finiteness of the first moment only, i.e., $\mathbb{E}\left(M_{u}^{*}\right)<\infty$, as the proof for the mgf is analogous.
$\mathbb{P}\left(M_{u}^{*}>k\right)$ $=\mathbb{P}\left[\sup\limits_{n\geq k}\bar{Z}_{n}>u\right]$
$\leq\mathbb{P}\left[\bigcup\limits_{n\geq k}\left(\bar{Z}_{n}\geq u\right)\right]$
$\leq\sum_{n=k}^{\infty}\mathbb{P}\left(\bar{X}_{n}\geq 0\right)\left[\text{since }\bar{X}_{n}=\bar{Z}_{n}-u\right]$
[…]
$=\frac{C\cdot\rho^{k}}{1-\rho}\quad[\text{this holds for large }k]$
Hence, for large $k$, $\mathbb{E}\left(M_{u}^{*}\right)\leq k+\frac{C}{(1-\rho)^{2}}\rho^{k}<\infty$. ∎
5 Simulation
In this section, we present simulation results to examine the limiting behavior of $N_{1,N}$ as the total sample size $N$ increases and to support our main result in Section 3. We implement the allocation rule described in Section 2 for various values of $(\theta_{0},\theta_{1},\sigma_{0},\sigma_{1})$ in the context of Normal populations. For each $N$, we conduct $10,000$ replications and use the simulated data to compute the probability of correct selection (PCS) as the proportion of correct decisions in these trials. In addition, we estimate $\mathbb{E}(N_{1,N})$ for different parameter settings and analyze its limiting values as $N$ increases.
Table 1: Simulation results for distinct Normal mean pairs $(\theta_{0},\theta_{1})$ and variance pairs $(\sigma_{0}^{2},\sigma_{1}^{2})=(1,0.7)$ under a fixed initial sample size.
Total Sample $(\theta_{0},\theta_{1})=(0.5,0)$ $(\theta_{0},\theta_{1})=(0.8,0.2)$ $(\theta_{0},\theta_{1})=(1,0.5)$
     Size $(N)$ PCS $\mathbb{E}(N_{1,N})$ PCS $\mathbb{E}(N_{1,N})$ PCS $\mathbb{E}(N_{1,N})$
        200 0.9396 16.4209 0.9718 15.7146 0.9363 16.4576
        300 0.9407 16.7092 0.9710 15.8912 0.9377 16.7389
        400 0.9412 16.8754 0.9709 15.9557 0.9415 16.8625
        800 0.9413 17.6627 0.9726 16.0868 0.9430 17.4870
        900 0.9467 17.1830 0.9735 16.1228 0.9429 17.7404
      1000 0.9421 17.5419 0.9732 16.2150 0.9405 17.2427
      1500 0.9428 18.3047 0.9679 16.3760 0.9407 18.1658
      2000 0.9452 18.1094 0.9717 16.4335 0.9426 18.1526
      2500 0.9452 18.1166 0.9718 16.4929 0.9432 18.0036
      3000 0.9409 18.4554 0.9715 16.5518 0.9344 18.3848
      3500 0.9423 18.5810 0.9722 16.5738 0.9376 18.9482
The simulation results for different mean pairs $(\theta_{0},\theta_{1})$ and initial sample sizes are summarized in Table 1. We simulate PCS values and $\mathbb{E}(N_{1,N})$ for several choices of $N$, ranging from moderate to large, across three sets of parameter values $(\theta_{0},\sigma_{0}^{2})$ and $(\theta_{1},\sigma_{1}^{2})$. The analysis indicates that $\mathbb{E}(N_{1,N})$ stabilizes to a constant as $N$ increases. Furthermore, we observe that the allocation rule assigns more samples to the population with the higher mean value. Each simulation starts with equal initial sample sizes for both populations, carefully chosen on the basis of the parameter settings.
Table 2: Simulation results for distinct Bernoulli probability pairs $(p_{0},p_{1})$ under a fixed initial sample size.
Total Sample $(p_{0},p_{1})=(0.5,0.2)$ $(p_{0},p_{1})=(0.6,0.3)$ $(p_{0},p_{1})=(0.8,0.5)$
     Size $(N)$ PCS $\mathbb{E}(N_{1,N})$ PCS $\mathbb{E}(N_{1,N})$ PCS $\mathbb{E}(N_{1,N})$
        200 0.9714 15.7156 0.9613 15.8792 0.9700 15.6926
        300 0.9704 15.8346 0.9623 15.9476 0.9667 15.7851
        400 0.9678 15.9646 0.9626 16.0014 0.9721 15.7609
        800 0.9690 16.0128 0.9648 16.5692 0.9732 15.9030
        900 0.9710 16.1097 0.9649 16.4865 0.9701 15.9566
      1000 0.9676 16.3659 0.9617 16.4393 0.9722 16.1128
      1500 0.9684 16.3968 0.9648 16.6400 0.9710 16.1403
      2000 0.9727 16.2998 0.9642 16.5592 0.9709 16.3007
      2500 0.9731 16.3938 0.9661 16.5206 0.9714 16.4970
      3000 0.9676 17.0703 0.9661 16.6232 0.9698 16.9098
      3500 0.9709 16.6486 0.9673 16.3406 0.9671 16.3561
Chernoff’s theorem also holds for other distributions, including Bernoulli variables. In a similar manner, we conduct simulations for Bernoulli populations, performing $10,000$ replications for three specific sets of values of parameters $(p_{0},p_{1})$, where $p_{i}$ denotes the probability of success for $i=0,1$. The corresponding simulation results are presented in Table 2.
The findings indicate that $\mathbb{E}(N_{1,N})$ tends to a constant as $N$ grows large. For each set of parameter values, we begin with equal initial sample sizes for both populations, ensuring a balanced allocation at the start of the procedure.
6 Conclusion
We have shown in this study that with correct inference, the number of applications of the less effective drug can be kept finite. We considered a simple and natural procedure and have applied Chernoff’s Theorem for proving the results, as it is applicable to a broader range of response distributions, and hence it has meaningful practical implications.